Grade 12Algebra

A vehicle covers a distance of 43.2 km in 2.4 liters of petrol. How much distance will it cover in one liter of petrol?

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question, it is given that,
Total distance covered by vehicle in 2.4 litres of petrol = 43.2 kmThen,
Distance covered in 1 litre of petrol = 43.2 ÷ 2.4
= (432/10) ÷ (24/10)
= (432/10) × (10/24)
= (432 × 10)/ (10 × 24)
= (36 × 1)/ (1 × 2)
= (18 × 1)/ (1 × 1)
= 18 km
∴Total distance covered in 1 liter of petrol is 18 km

Thanks